OpenStudy (mathmath333):

Two fair dice are thrown. Find the probablity of getting a number divisible by 2 or 4

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ Two fair dice are thrown. Find the probablity}\hspace{.33em}\\~\\ & \normalsize \text{ of getting a number divisible by 2 or 4}\hspace{.33em}\\~\\ & a.)\ \dfrac{1}{2} \hspace{.33em}\\~\\ & b.)\ \dfrac{3}{4} \hspace{.33em}\\~\\ & c.)\ \dfrac{1}{3} \hspace{.33em}\\~\\ & d.)\ \dfrac{2}{3} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (jhulian):

The probability of a number divisible by 2 (2,4,6) => 3/6 =1/2 Probability of number divisible by 4 (4) => 1/6 Final probability => (1/2) / (1/6) => 2/3

OpenStudy (jhulian):

Final probability: W

OpenStudy (jhulian):

sorry...typos:-)

OpenStudy (jhulian):

Final probability: We have to add (1/2) + (1/6) = 2/3

OpenStudy (mathmath333):

but the answer given is a.) 1/2

Nnesha (nnesha):

`two dice ` not just one Jhulian :=))

OpenStudy (ikram002p):

2 dice, and at least one is divisible by 2 or 4 ? or exactly one?

OpenStudy (ikram002p):

or is it the sum of both ?

OpenStudy (mathmath333):

i m also confused about that sum or individual thing that y posted here :)

OpenStudy (ikram002p):

i see :) lets solve in either way

OpenStudy (anonymous):

The correct answer is 1/2 = A

Nnesha (nnesha):

i assumed sum should be divisible by 2 or 4 so \[(1,1)( 1,3)(1,5) (2,2)(2,4)(2,6)(3,3)(3,5)(4,4)(4,6)((5,5)(6,6)\] total = 36 12/36 are divisible by 2 \[\color{Red}{( 1,3)(2,2)(2,6)(3,5)(4,4)(6,6)}\] and 6/36 divisible by 4 \[\frac{ 12 }{ 36}+\color{ReD}{\frac{6}{36}}\]

OpenStudy (jhulian):

OMG...My bad :-)